OpenStudy (anonymous):

Four friends went scuba diving. According to the following statements, who spotted sea life at the fastest rate? A. Lance saw 2 reef sharks in 5 minutes. B.Jasmine saw 4 sea cucumbers in 8 minutes. C. Robert saw 5 jellyfish in fifteen minutes. D.Shilpa saw 3 pencil urchins in ten minutes.

OpenStudy (anonymous):

A. 2/5 = .4 B. 4/8 = .5 C. 5/15 = .33333 D. 3/10 = .3 So B.

OpenStudy (anonymous):

thanks